# Work Statistics via Real-Time Effective Field Theory: Application to Work Extraction from Thermal Bath with Qubit Coupling

### v2

@@ I Introduction
Before proceeding to the main text, we remark on some subtle points:
(I) Although the original consideration of the JC-FT and work statistics in Jarzynski (1997); Crooks (2000, 1999); Tasaki (2000) is only for the pure thermal states, it can be naturally extended to the diagonal states rather than the thermal states in the conventional two-point measurement (TPM) scheme, for example, the case as in Fig. 1(b). Especially, the WDF and WCF can still be defined. We have illustrated this extension explicitly in Appendix A. Moreover, the JC-FT (1) no longer holds and should be modified into Evans et al. (1993); Evans and Searles (1994); Gallavotti and Cohen (1995); Crooks (1999)
@@
(II) As mentioned, the setup Fig. 1(b) can be viewed as a heat engine/refrigerator with two heat baths at temperatures $T_{B}$ and $T_{Q}$, respectively. Assume $Q_{B}$ and $Q_{Q}$ are, respectively, the heat flows into the bath and the qubit. Then, by the first law relations of the adiabatic processes [^3] This is consistent with the EFT consideration by requiring the externally driven coupling for the cyclic processes to be adiabatically varying in time., one has
@@
(III) We will not introduce a work storage system, but will implicitly assume the external agent can carry out the task. This is similar to the original consideration in Jarzynski (1997); Crooks (2000, 1999); Tasaki (2000)  [^4] The external agent is either explicitly mentioned in Tasaki (2000), or implicitly introduced in Jarzynski (1997); Crooks (2000, 1999) by emphasizing the work is done on the bath by evolving the external parameters. , and see Alhambra et al. (2016); Åberg (2018) for the consideration with an additional storage system.
(IV) If the thermal stats are replaced by non-diagonal ones, the TPM will then destroy the initial coherence, so that one cannot define the WDF or WCF to satisfy some general fluctuation theorem. Instead, the entire driven process should be treated as a quantum channel, and the reverse process for the generalized fluctuation theorem is the Petz recovery map Petz (1986), with the initial non-diagonal state serving as the reference state. The generalized fluctuation theorem will then be reformulated as constraints on the exchange of information and entropy production Alhambra et al. (2016); Åberg (2018); Kwon and Kim (2019b); Bai et al. (2024). This, however, falls beyond the scope of our current study.
The rest of the paper is organized as follows. In the next section, we provide a brief review of work statistics in the real-time formalism and extend it to the EFT context. In section III, we apply the EFT formalism to obtain the work statistics of the pure thermal bath, and then in section IV to the cases with qubit couplings to the thermal bath. In section V, we present the numerical patterns of work extraction based on the analytical results in section IV. We then conclude this work in section VI. Technical details are relegated to the Appendices. We give a thorough review in Appendix A on the real-time formalism of work statistics. In Appendix B, we review the generalized Wick’s theorem and then apply it to derive the non-perturbative work statistics for pure thermal states. In Appendix C and D, we give the details of deriving work statistics for the spin qubit coupling and fermion/topological qubit coupling, respectively. Finally, we provide some supplementary plots in Appendix E.
II Work statistics via real-time EFT
@@
We use the relations (II) to rewrite (18) into the following
@@
III Work statistics of pure thermal bath
@@
IV Work statistics of thermal bath with qubit coupling
@@
V Work extraction by qubit coupling and its efficiency
@@
VI Conclusion
In this paper, we have adopted the effective field theory approach to calculate the work statistics of an irreversible process without and with qubit coupling to the thermal bath. With such an efficient tool, we can explicitly calculate the work characteristic function, work distribution function, and possible work extraction by varying the model parameters, such as the type of Ohmic spectral functions of the quasiparticles, temperature of the thermal bath, and the spectral gap of the qubit system. Especially with our method, we can derive the nonperturbative work distribution for the thermal states. As far as we know, this is the first nonperturbative result obtained in the literature. We also find that the work extractable regime by the qubit coupling is usually restricted but can be pinned down precisely from the results based on our framework. This will be useful for designing the quantum engine. Furthermore, we can treat the qubit-bath system as a heat engine or refrigerator, and derive the corresponding efficiency or COP from work statistics by assuming first law relations of adiabatic processes. Our results indicate the spin qubit coupling yields a far efficient heat engine than the other two alternatives, but it is the topological qubit coupling when wanting a far better refrigerator. This implies that the underlying quantum statistics of the qubit-bath system are quite relevant for designing a quantum heat engine or refrigerator.
@@ VI Conclusion
Acknowledgements. This work is supported by Taiwan’s National Science and Technology Council (NSTC) through Grant No. 112-2112-M-003-006-MY3. We thank the referee of PRE for the inspiring suggestion of considering the efficiency of a quantum engine using the first law relations.

### v3

@@ I Introduction
The rest of the paper is organized as follows. In the next section, we extend the fluctuation theorem for the work statistics to that of entropy production for non-thermal states, and derive the efficiency/coefficient of performance of the heat engine/refrigerator for the setup Fig. 1(b). In section III, we provide a brief review of work statistics in the real-time formalism and extend it to the EFT context. In section IV, we apply the EFT formalism to obtain the work statistics of the pure thermal bath, and then in section V to the cases with qubit couplings to the thermal bath. In section VI, we present the numerical patterns of work extraction based on the analytical results in section V. We then conclude this work in section VII. Technical details are relegated to the Appendices. We give a thorough review in Appendix A on the real-time formalism of work statistics. In Appendix B, we review the generalized Wick’s theorem and then apply it to derive the non-perturbative work statistics for pure thermal states. In Appendix C and D, we give the details of deriving work statistics for the spin qubit coupling and fermion/topological qubit coupling, respectively. Finally, we provide some supplementary plots in Appendix E.
II Entropy production, work statistics, heat engine/refrigerator and all that
Although the original consideration of the JC-FT and work statistics in Jarzynski (1997); Crooks (2000, 1999); Tasaki (2000) is only for the pure thermal states, it can be naturally extended to the diagonal states rather than the thermal states in the conventional two-point measurement (TPM) scheme, for example, the case as in Fig. 1(b). Especially, the WDF and WCF can still be defined. We have illustrated this extension explicitly in Appendix A. Moreover, the JC-FT (1) no longer holds and should be modified into Evans et al. (1993); Evans and Searles (1994); Gallavotti and Cohen (1995); Crooks (1999)
@@ II Entropy production, work statistics, heat engine/refrigerator and all that
As mentioned, the setup Fig. 1(b) can be viewed as a heat engine/refrigerator with two heat baths at temperatures $T_{B}$ and $T_{Q}$, respectively. Assume $Q_{B}$ and $Q_{Q}$ are, respectively, the heat flows into the bath and the qubit. Then, by the first law relations of the adiabatic processes [^3] This is consistent with the EFT consideration by requiring the externally driven coupling for the cyclic processes to be adiabatically varying in time., one has
@@
We will not introduce a work storage system, but will implicitly assume the external agent can carry out the task. This is similar to the original consideration in Jarzynski (1997); Crooks (2000, 1999); Tasaki (2000)  [^4] The external agent is either explicitly mentioned in Tasaki (2000), or implicitly introduced in Jarzynski (1997); Crooks (2000, 1999) by emphasizing the work is done on the bath by evolving the external parameters. , and see Alhambra et al. (2016); Åberg (2018) for the consideration with an additional storage system. We explicitly calculate in section VI the efficiency/COP of the heat engine/refrigerator of Fig. 1(b) for the qubit-bath coupling in the framework of real-time EFT.
If the thermal stats are replaced by non-diagonal ones, the TPM will then destroy the initial coherence, so that one cannot define the WDF or WCF to satisfy some general fluctuation theorem. Instead, the entire driven process should be treated as a quantum channel, and the reverse process for the generalized fluctuation theorem is the Petz recovery map Petz (1986), with the initial non-diagonal state serving as the reference state. The generalized fluctuation theorem will then be reformulated as constraints on the exchange of information and entropy production Alhambra et al. (2016); Åberg (2018); Kwon and Kim (2019b); Bai et al. (2024). This, however, falls beyond the scope of our current study.
III Work statistics via real-time EFT
@@
We use the relations (III) to rewrite (18) into the following
@@
IV Work statistics of pure thermal bath
@@
V Work statistics of thermal bath with qubit coupling
@@
VI Work extraction by qubit coupling and its efficiency
@@
VII Conclusion
In this paper, we have adopted the effective field theory approach to calculate the work statistics of an irreversible process without and with qubit coupling to the thermal bath. With such an efficient tool, we can explicitly calculate the work characteristic function, work distribution function, and possible work extraction by varying the model parameters, such as the type of Ohmic spectral functions of the quasiparticles, temperature of the thermal bath, and the spectral gap of the qubit system. Especially with our method, we can derive the nonperturbative work distribution for the thermal states. As far as we know, this is the first nonperturbative result obtained in the literature. We also find that the work extractable regime by the qubit coupling is usually restricted but can be pinned down precisely from the results based on our framework. This will be useful for designing the quantum engine. Furthermore, we can treat the qubit-bath system as a heat engine or refrigerator, and derive the corresponding efficiency or COP from work statistics by assuming first law relations of adiabatic processes. Our results indicate the spin qubit coupling yields a far efficient heat engine than the other two alternatives, but it is the topological qubit coupling that yields a far better refrigerator. This implies that the underlying quantum statistics of the qubit-bath system are quite relevant for designing a quantum heat engine or refrigerator.
@@ VII Conclusion
Acknowledgements. This work is supported by Taiwan’s National Science and Technology Council (NSTC) through Grant No. 112-2112-M-003-006-MY3. We thank the anonymous referee of PRE for the inspiring suggestion of considering the efficiency of a quantum engine using the first law relations.